OpenStudy (anonymous):

Divide a+7/a-5 divided by a-4/a+8

OpenStudy (anonymous):

(a+7)(a+8)/(a-5)(a-4)

OpenStudy (anonymous):

((7+a) (8+a))/((-5+a) (-4+a)) = (7 (8+a)+a (8+a))/((-5+a) (-4+a)) = (7 (8+a)+a (8+a))/((-5+a) (-4+a)) = (7 (8+a)+a (8+a))/(-5 (-4+a)+a (-4+a)) = (7 (8+a)+a (8+a))/(-5 (-4+a)+a (-4+a)) = (7 (8+a)+a 8+a a)/(-5 (-4+a)+a (-4)+a a) = (7 (8+a)+8 a+a a)/(-5 (-4+a)-4 a+a a) = (7 8+7 a+8 a+a a)/(-5 (-4)-5 a-4 a+a a) = (56+7 a+8 a+a a)/(20-5 a-4 a+a a) = (56+7 a+8 a+a^2)/(20-5 a-4 a+a^2) = (a^2+15 a+56)/(a^2-9 a+20)